OpenStudy (anonymous):

What is the solution to the following system of equations? y = -x2 - 5x - 4 y = -x2 + 9x - 18 (2 points) (-1, -10) (1, -10) (-1, 10) (1, 10)

OpenStudy (anonymous):

(-1, 10) is that i got im in a super hurry though i have to go to bed in like 10 mins

OpenStudy (anonymous):

what i got*

OpenStudy (marigirl):

no i got (1, -10)

OpenStudy (anonymous):

Can you show me how i think i did my math wrong

OpenStudy (anonymous):

oh wait i think i flipped the negative

OpenStudy (whpalmer4):

so if you plug x = -1 into both equations, do you get y = 10? if so, that's a solution. if not, it isn't. y = -x^2 - 5x - 4 10 = -(-1)^2 - 5(-1) -4 10 = -1 + 6 - 4 10 = 1 oops. not a solution! let's try @marigirl's answer y = -x^2 - 5x - 4 -10 = -(1)^2 - 5(1) - 4 -10 = -1 - 5 - 4 -10 = -10 good so far, but it has to satisfy all the equations in the system y = -x^2 + 9x - 18 -10 = -(1)^2 + 9(1) - 18 -10 = -1 + 9 - 18 -10 = -10 (1,-10) is the solution to the system of equations.

OpenStudy (whpalmer4):

remember, a common mistake is thinking that -x^2 = (-x)*(-x). it doesn't. it's -1*x*x.

OpenStudy (marigirl):

-x^2 - 5x - 4 -x^2 + 9x - 18 - minus one from another --------------- 0x^2 -14x+14 now solve for x -14x=-14 x=1 then plug in x=1 into any of the original equations. (1, -10)

OpenStudy (anonymous):

thank you very much i see my mistake

OpenStudy (whpalmer4):

as for how to solve it, at the solution, obviously y will have the same value for both equations, right? so we can say -x^2 - 5x -4 = -x^2 + 9x - 18 add x^2 to both sides -5x -4 = 9x - 18 add 5x to both sides -4 = 14x - 18 add 18 to both sides 14 = 14x x = 1

OpenStudy (whpalmer4):

you can also do it @marigirl's way, but be very careful about the signs when subtracting polynomials. I prefer to multiply one polynomial by -1, then add the two.